# SPADE: Spatial Transcriptomics and Pathology Alignment Using a Mixture of Data Experts for an Expressive Latent Space

### v1

@@ Abstract
The rapid growth of digital pathology and advances in self-supervised deep learning have enabled the development of foundational models for various pathology tasks across diverse diseases. While multimodal approaches integrating diverse data sources have emerged, a critical gap remains in the comprehensive integration of whole-slide images (WSIs) with spatial transcriptomics (ST), which is crucial for capturing critical molecular heterogeneity beyond standard hematoxylin & eosin (H&E) staining. We introduce SPADE, a foundation model that integrates histopathology with ST data to guide image representation learning within a unified framework, in effect creating an ST-informed latent space. SPADE leverages a mixture-of-data experts technique, where experts, created via two-stage feature-space clustering, use contrastive learning to learn representations of co-registered WSI patches and gene expression profiles. Pre-trained on the comprehensive HEST-1k dataset, SPADE is evaluated on 14 downstream tasks, demonstrating significantly superior few-shot performance compared to baseline models, highlighting the benefits of integrating morphological and molecular information into one latent space.
* Indicates shared first authorship
@@ 1 Introduction
High-resolution whole slide images (WSIs) have propelled the development of powerful deep learning foundation models in computational pathology, demonstrating robust performance across diverse tissue types and tasks [29, 15, 25, 10]. These models excel in various applications, from cancer subtype classification to patient outcome prediction.
Due to their gigapixel resolution and limited pixel-level annotations, WSIs are typically processed as patches. Standard practice is to use multiple instance learning (MIL), which utilizes weak slide-level labels [4] and involves embedding individual patches with a pre-trained model and then aggregating these embeddings for downstream tasks. This approach relies on high-quality patch embeddings, driving the development of self-supervised foundation models in digital pathology [28, 6, 5, 32]. When applied to images or paired data from multiple modalities, self-supervised learning produces diverse, expressive feature encoders and enables training on large unlabeled datasets spanning multiple organs. These task-agnostic self-supervised encoders are robust to institutional variations, including differences in staining procedures and other image-quality factors. [28, 6, 5, 32].
@@
SPADE’s novel architecture based on a mixture-of-experts is specifically designed to handle the complexities of spatially resolved paired ST and H&E data. Each expert is trained on a distinct subset of ST-WSI pairs, identified through a two-step K-means clustering process. First, patches are grouped into K subclusters within each organ, yielding per-organ centroids. Then, these centroids are clustered again across all organs to arrive at a final set of K representative clusters spanning the 515 Visium WSIs. This process generates hard negatives, enhancing training for individual experts and ensuring robust representation learning across multiple tissue types. Our contributions are summarized as follows:
@@
    Mixture-of-Experts with Hard Negative Mining: By clustering WSIs in a two-step K-means process, SPADE identifies challenging examples that enhance discrimination and prevent trivial solutions, especially across diverse organ types.
@@
    Comprehensive Evaluation across 14 Tasks: We demonstrate SPADE’s effectiveness on eight cancer classification tasks and six survival prediction tasks, showcasing its robust performance in modeling the complexity of tumor samples.
@@ 2.1 Contrastive learning pretraining for multimodal datasets
Figure 2: Interpretability of SPADE. Attention scores for a prostate WSI positive for biochemical recurrence (BCR) are visualized as a heatmap for the UNI and SPADE models. The deeper the red color the higher attention the model put on that region of the tissue. The yellow outline shows the predicted tumor boundary. SPADE shows higher concentration of attention within the tumor boundary. Sub-crops are selected to show the patterns with high-attention at higher magnifications.
BLEEP applies the CLIP framework to spatial transcriptomics for gene expression prediction, creating a multimodal model specifically for Visium data but trained on a single tissue type, limiting its generalizability. Another model, TANGLE, recently developed by Jaume et al. [13] applies the CLIP framework to paired H&E WSIs and bulk RNA-seq data from the Cancer Genome Atlas (TCGA); however, bulk RNA sequencing requires tissue dissociation and lacks the spatial resolution present in each WSI. SPADE addresses these limitations by training on multiple tissue types using a Mixture-Of-Experts approach, utilizing paired H&E and ST data at the patch level to learn rich spatially-informed patch-level embeddings for multiple tissue types.
@@ 2.2 Clustering for non-trivial contrastive learning
SPADE is trained on a dataset that spans multiple organs leading to a risk of the original CLIP framework finding trivial solutions by contrasting different tissue types with each other instead of learning a meaningful latent space. Because CLIP samples negative pairings from within each batch, highly varied images in a batch can make these pairings easier for the model to learn. Ma et al. proposed a Mixture of Data Experts (MoDE) approach to address the limitations of CLIP using non-trivial negative pairings to train by incorporating a two-step K-means clustering process into training [19]. MoDE first clusters semantically similar text captions, then trains a distinct CLIP model within each cluster. By sampling hard negatives from within these similar groups, the model is forced to focus on distinguishing features that differentiate similar captions and their paired images [19].
Using a two-step K-means procedure, we adapt this hard-negative sampling procedure to Visium data and cluster in the image feature space. In [27] the authors demonstrate how hierarchical clustering can be leveraged to create high-quality representations by progressively refining clusters at different levels of abstraction. Inspired by this, our two-step K-means procedure first clusters data within each organ type, and then clusters the resulting organ-specific centroids. This allows our model to learn a more detailed and expressive latent space for each expert while also allowing flexibility to work across cancer types and vastly different tissue makeups.
3 Method
We present our framework: SPADE, Spatial transcriptomics and Pathology Alignment using a mixture of Data Experts. (Figure 1). The proposed framework acts as an image encoder, where the extracted embeddings are subsequently aggregated to generate a comprehensive slide-level representation. The proposed framework comprises a mixture of data experts $\{f(\cdot|c)\}$ (Section 3.3), where each data expert $f(\cdot|c)$ is a CLIP-based model trained on the cluster $c$ (Section 3.2) using a contrastive objective to align an H&E patch embedding with a gene expression embedding (Section 3.1). SPADE is tested on a variety of downstream tasks (Section 4.4). Additional details can be found in Supplementary Materials.
3.1 Background: Basic contrastive pretraining
SPADE utilizes a contrastive learning framework inspired by CLIP, designed with a modified loss function to learn a bimodal embedding that aligns image patches with their corresponding gene expression data [31]. In contrast to natural images, it is common to sample multiple spots with similar gene expression profiles or image morphological features within the same batch, which the contrastive objective should not pull apart. To take into account similar pairs, the similarity-adjusted target $t$ is computed as follows:
@@ 3.1 Background: Basic contrastive pretraining
where $h_{v}$ is an $m$-dimensional image embedding, $h_{x}$ is an $m$-dimensional gene expression embedding, and $\tau$ temperature parameter. The contrastive objective $\mathcal{L}$ is then updated to align $h_{v}$ and $h_{x}$ in the latent space and is computed with cross-entropy (CE) as follows:
@@ 3.2 Clustering
Each WSI $X_{j},j=1,\dots,J$ from the pretraining cohort $\mathcal{D}$ of size $J$ is divided into a set of non-overlapping patches $X_{j}=\{x_{j}^{1},\dots,x_{j}^{N_{j}}\}$, where each patch $x_{j}^{n}\in\mathbb{R}^{W\times H\times 3}$. A feature encoder $f_{\text{enc}}(\cdot)$ pre-trained on a large cohort of histopathology data is then used to extract a compressed embedding from each patch. This results in a set of embeddings $H_{v,j}=\{h_{v,j}^{1},\dots,h_{v,j}^{N_{j}}\}$ where $h_{v,j}^{n}=f_{\text{enc}}(x_{j}^{n})\in\mathbb{R}^{m}$.
The goal is to extract a set of image clusters $\{h_{v,1},...,h_{v,C}\}$ with $h_{v,c}\in\mathbb{R}^{m}$ based on the set of patch-level embeddings to formulate conditions $C$ for data expert training. To train these experts, we utilize a two-step K-means clustering approach. In the first step, we apply fine-grained clustering utilizing patch-based embeddings within each organ type to identify clusters where the samples share similar semantics. In the second step, we aggregate the fine-grained clusters based on each cluster’s centroid to establish coarse-grained clustering, facilitating specialization among data experts across different organs.
@@
, where $S=\bigcup S_{i}$ set of learned cluster centers.
@@
To determine the optimal number of clusters for both steps, we performed a within-cluster sum of squares (WCSS) analysis and used the elbow method to identify the point where adding more clusters provided diminishing returns. This ensured that our clustering approach captured meaningful biological variation while avoiding excessive fragmentation of the data
Lung Prostate
Lymph
Nodes
Ovarian Breast
PLCO CPTAC PANDA
TCGA-
PRAD
CAM16
UBC-
OCEAN
PLCO
TCGA-
BRCA
AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC F1
TANGLE 95.9 91.1 96.7 96.6 88.1 57.4 - - 74.2 57.9 96.4 81.5 76.4 55.3 - -
GigaPath 97.0 90.1 97.1 89.9 72.5 33.9 75.3 38.7 93.0 88.3 94.2 77.9 69.6 50.4 69.4 63.9
UNI+ABMIL 97.1 89.3 96.9 88.8 93.2 66.8 81.0 52.7 96.0 81.7 97.8 89.1 79.7 55.0 97.5 97.1
CONCH+ABMIL 97.0 91.3 96.7 88.8 90.9 62.3 78.9 47.9 81.8 74.4 96.5 82.9 76.7 51.6 98.3 96.7
SPADE+ABMIL 98.5 87.1 97.9 92.5 94.9 74.7 82.1 56.9 98.6 89.1 97.8 87.1 82.1 57.3 97.7 98.9
UNI+TransMIL 97.0 77.8 93.1 82.1 92.1 64.8 76.2 47.5 96.9 85.7 97.1 82.0 70.4 47.3 97.8 97.3
CONCH+TransMIL 98.3 89.3 98.1 91.3 92.1 63.8 80.9 46.8 96.3 88.3 95.7 83.6 76.9 53.8 97.2 95.9
SPADE+TransMIL 96.6 82.0 94.6 89.7 91.7 62.5 72.0 43.9 93.4 82.9 97.3 83.1 72.5 47.9 97.3 97.5
@@ 3.3 Data Experts Training
In this work, we propose to train a set of CLIP models $\{f(\cdot|c)\}$ independently given clusters $C$, such that $c\in C$. The training data for a data expert $\{f(\cdot|c)\}$ is :
@@ 3.4 Slide-level Representation
The proposed experts-based approach provides multiple patch-based models to select from during inference rather than relying solely on a single CLIP model. We evaluate three distinct weighting techniques for routing data experts by formulating their outputs as a weighted sum for simplicity:
@@
, where $d_{c^{\prime}}=(||h_{v,s}-h_{v,c^{\prime}})||^{2}_{2}$. In the second approach, the outputs from all experts are averaged for each data sample:
@@
In the third approach, the outputs of all experts are combined using a normalized weighted sum, with the weights being inversely proportional to the distance from the data sample to each expert’s centroids:
$p_{3}(c|(h_{v,s},h_{v,c}))=\frac{\bar{d_{c}}}{\sum_{C}\bar{d_{i}}},\bar{d_{c}}=\frac{1}{d_{c}}$ (9)
Prostate BCR Lung Uterine Breast
Private TCGA-PRAD PLCO TCGA-UCEC TCGA-BRCA PLCO
c-index c-index c-index c-index c-index c-index
TANGLE 60.3 - 48.6 - - 57.0
GigaPath 66.4 53.0 59.5 69.1 65.5 56.0
UNI+ABMIL 58.8 56.5 57.8 64.6 74.1 56.0
CONCH+ABMIL 60.3 55.9 35.8 64.5 64.5 55.5
SPADE+ABMIL 67.0 59.4 64.7 72.4 71.8 59.5
UNI+TransMIL 56.5 57.7 58.4 72.2 70.3 54.2
CONCH+TransMIL 52.5 53.0 46.8 65.9 65.9 54.7
SPADE+TransMIL 71.4 58.8 72.3 71.0 73.0 57.4
Table 2: Survival and BCR prediction results of SPADE and baselines for six different subtyping tasks. c-index values are shown as percentages. The best performance is in bold, and the second best is underlined. TANGLE could not be applied to TCGA datasets since they were used to train the model.
After features are obtained for each patch, a slide-level feature vector can be derived through a learnable aggregation method. Specifically, we employed Attention-Based Multiple Instance Learning (ABMIL) [11] and Transformer-based Multiple Instance Learning (TransMIL) [26], which allow for aggregating patch-level embeddings into a slide-level representation. In this setup, each patch embedding is treated as an instance, and the model learns to assign attention weights to each instance based on its relevance to the slide-level prediction task. ABMIL leverages an attention mechanism to weight each patch’s contribution, capturing spatial heterogeneity across the slide. In contrast, TransMIL uses transformer-based self-attention layers to capture dependencies between patches, allowing for more complex interactions and relationships to be modeled. The resulting slide-level feature vector can then be utilized for various downstream tasks, as described in Section 4.1.
@@ 4.1 Datasets
Lung Prostate Lymph Nodes Ovarian Breast
PLCO CPTAC PANDA CAM16 UBC-OCEAN PLCO
AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC F1
$p_{1}$+ABMIL 97.9 85.0 97.1 91.1 94.4 71.6 97.1 87.5 96.8 76.9 79.3 51.4
$p_{2}$+ABMIL 97.6 95.2 97.7 90.4 94.0 711 93.6 77,6 97.2 80.0 76.6 55.6
$p_{3}$+ABMIL 98.5 87.1 97.9 92.5 94.9 74.7 98.6 89.1 97.8 87.1 82.1 57.3
$p_{1}$+TransMIL 96.5 82.0 95.4 86.0 88.9 55.7 77.7 67.5 95.0 75.9 70.7 47.6
$p_{2}$+TransMIL 91.4 70.0 94.9 85.7 90.2 59.9 92.9 83.0 96.4 82.9 77.8 54.0
$p_{3}$+TransMIL 97.4 93.0 96.3 89.7 91.7 62.5 93.4 82.9 97.3 83.1 72.5 47.9
Each dataset was processed in a similar manner with a fixed stain normalization and 224$\times$224 patching at magnification 20x. The Visium gene expression data was processed by normalizing each spot to its total count, followed by log normalization. Subsequently, highly variable genes were identified using the Scanpy package [30]. The union of the top 50 most highly variable genes from each sample was used for training and prediction, resulting in 7,986 genes following the method outlined in [31]. Additional details can be found in Supplementary Materials.
@@
Subtyping We evaluated the proposed approach on six different subtyping tasks: Non-Small Cell Lung Carcinoma (NSCLC) subtyping on PLCO [34] and CPTAC [9] (two classes), ISUP grading based on Prostate cancer grade assessment (PANDA) challenge [3] (six classes), classification of breast cancer metastases in lymph nodes (Camelyon16) [2, 16] (three classes), ovarian cancer subtypes classification (UBC-OCEAN) challenge [8, 1] (five classes), and PLCO Breast cancer subtyping classification (three classes) [34]. For each dataset where patient identifiers were provided, all patient WSIs were aggregated into one set of vectors for input to ABMIL and TransMIL.
Survival We evaluated the proposed method on four different survival tasks: Breast Invasive Carcinoma (BRCA) from TCGA and PLCO, Non-Small Cell Lung Carcinoma (NSCLC) from PLCO, and Uterine Corpus Endometrial Carcinoma (UCEC) from TCGA.
Prostate Cancer Biochemical Recurrence We evaluated the proposed solution on two different prostate cancer datasets, one public coming from TCGA (TCGA-PRAD) and one private dataset coming from our institution.

### v2

@@ Abstract
The rapid growth of digital pathology and advances in self-supervised deep learning have enabled the development of foundational models for various pathology tasks across diverse diseases. While multimodal approaches integrating diverse data sources have emerged, a critical gap remains in the comprehensive integration of whole-slide images (WSIs) with spatial transcriptomics (ST), which is crucial for capturing critical molecular heterogeneity beyond standard hematoxylin & eosin (H&E) staining. We introduce SPADE, a foundation model that integrates histopathology with ST data to guide image representation learning within a unified framework, in effect creating an ST-informed latent space. SPADE leverages a mixture-of-data experts technique, where experts are created via two-stage imaging feature-space clustering using contrastive learning to learn representations of co-registered WSI patches and gene expression profiles. Pre-trained on the comprehensive HEST-1k dataset, SPADE is evaluated on 20 downstream tasks, demonstrating significantly superior few-shot performance compared to baseline models, highlighting the benefits of integrating morphological and molecular information into one latent space. Code and pretrained weights are available at https://github.com/uclabair/SPADE.
journal: Medical Image Analysis
[…]
[label1]organization=Biomedical AI Research Lab, University of California, Los Angeles,addressline=24 Westwood Blvd, city=Los Angeles, postcode=90024, state=CA, country=United States
[…]
[label2]organization=Department of Pathology, University of California, Los Angeles,city=Los Angeles, state=CA, country=United States
[…]
[label3]organization=Department of Radiology, University of California, Los Angeles, city=Los Angeles, state=CA, country=United States
[…]
[label4]organization=Department of Bioengineering, University of California, Los Angeles, city=Los Angeles, state=CA, country=United States
[…]
[label5]organization=UCLA Medical Informatics Home Area, University of California, Los Angeles, city=Los Angeles, state=CA, country=United States
[…]
[label6]organization=Department of Computational Medicine, University of California, Los Angeles, city=Los Angeles, state=CA, country=United States
@@ 1 Introduction
High-resolution whole slide images (WSIs) have propelled the development of powerful deep learning foundation models in computational pathology, demonstrating robust performance across diverse tissue types and tasks [1, 2, 3, 4]. These models are typically trained using self-supervision, enabling learning from large unlabeled datasets and producing embeddings robust to institutional variations, including differences in staining procedures and other image-quality factors [5, 6, 7, 8].
@@
SPADE’s architecture, based on a mixture-of-experts, is specifically designed to handle the complexities of spatially resolved paired ST and H&E data. Each expert is trained on a distinct subset of ST-WSI pairs, identified through a two-step K-means clustering process. First, patches are grouped into K subclusters within each organ, yielding per-organ centroids. Then, these centroids are clustered again across all organs to arrive at a final set of K representative clusters spanning the 515 Visium WSIs. This process generates hard negatives, enhancing training for individual experts and ensuring robust representation learning across multiple tissue types. Our contributions are summarized as follows:
@@
    Mixture of Data Experts with Hard Negative Mining: By clustering WSIs in a two-step K-means process, SPADE identifies challenging examples that enhance discrimination and prevent trivial solutions, especially across diverse organ types.
@@
    Comprehensive Evaluation across 20 Tasks: We demonstrate SPADE’s effectiveness on eight cancer classification tasks, six survival prediction tasks, and six biomarker prediction tasks, showcasing its robust performance in modeling the complexity of tumor samples.
@@ 2.1 Contrastive learning pretraining for multimodal datasets
[…]
@@ 2.2 Clustering for non-trivial contrastive learning
SPADE is trained on a dataset that spans multiple organs, leading to a risk of the original CLIP framework finding trivial solutions by contrasting different tissue types with each other instead of learning a meaningful latent space. Because CLIP samples negative pairings from within each batch, highly varied images in a batch can make these pairings easier for the model to learn. Ma et al. proposed a Mixture of Data Experts (MoDE) approach to address the limitations of CLIP using non-trivial negative pairings to train by incorporating a two-step K-means clustering process into training [19]. MoDE first clusters semantically similar text captions, then trains a distinct CLIP model within each cluster. By sampling hard negatives from within these similar groups, the model is forced to focus on distinguishing features that differentiate similar captions and their paired images [19].
Using a two-step K-means procedure, we adapt this hard-negative sampling procedure to Visium data and cluster in the image feature space. In [20] the authors demonstrate how hierarchical clustering can be leveraged to create high-quality representations by progressively refining clusters at different levels of abstraction. Inspired by this, our two-step K-means procedure first clusters data within each organ type, and then clusters the resulting organ-specific centroids. This clustering approach allows our model to learn a more detailed and expressive latent space for each expert while also allowing flexibility to work across cancer types and vastly different tissue makeups.
3 Methods
We present the Spatial transcriptomics and Pathology Alignment using a mixture of Data Experts (SPADE) framework. (Figure 1). The proposed framework acts as an image encoder, where the extracted embeddings are subsequently aggregated to generate a comprehensive slide-level representation for downstream applications. SPADE is a mixture of data experts $\{f(\cdot|c)\}$, each pretrained using a contrastive objective to align an H&E patch embedding with a gene expression embedding (Section 3.1) derived from a specific subset of data $c$ as defined in Section 3.2. Following the training of data experts (Section 3.3), SPADE is comprehensively evaluated on diverse downstream tasks to validate its effectiveness (Section 4.4). Implementation details can be found in B.
3.1 Multimodal alignment
SPADE utilizes a contrastive learning framework inspired by CLIP, designed with a modified loss function to learn a bimodal embedding that aligns image patches with their corresponding gene expression data [17]. In contrast to natural images, it is common to sample multiple spots with similar gene expression profiles or image morphology within the same batch, which the contrastive objective should not pull apart. To take into account similar pairs, the similarity-adjusted target $t$ is computed as follows:
@@ 3.1 Multimodal alignment
where $h_{v}$ is an $m$-dimensional image embedding, $h_{x}$ is an $m$-dimensional gene expression embedding, and $\tau$ is a temperature parameter. The contrastive objective $\mathcal{L}$ is then updated to align $h_{v}$ and $h_{x}$ in the latent space and is computed with cross-entropy (CE) as follows:
@@ 3.2 Clustering
As shown in Figure 2, each WSI $X_{j},j=1,\dots,J$ from the pretraining cohort $\mathcal{D}$ of size $J$ is first divided into a set of non-overlapping patches $X_{j}=\{x_{j}^{1},\dots,x_{j}^{N_{j}}\}$, where each patch $x_{j}^{n}\in\mathbb{R}^{W\times H\times 3}$. A pre-trained foundation model (UNI [6] ) serves as the feature encoder $f_{\text{enc}}(\cdot)$ used to extract a compressed embedding from each patch. This results in a set of embeddings $H_{v,j}=\{h_{v,j}^{1},\dots,h_{v,j}^{N_{j}}\}$ where $h_{v,j}^{n}=f_{\text{enc}}(x_{j}^{n})\in\mathbb{R}^{m}$.
The goal is to extract a set of image clusters $\{h_{v,1},...,h_{v,C}\}$ with $h_{v,c}\in\mathbb{R}^{m}$ based on the set of patch-level embeddings to formulate conditions $C$ for data expert training. To train these experts, we utilize a two-step K-means clustering approach (see Figure 2). In the first step, we apply fine-grained clustering utilizing patch-based embeddings within each organ type to identify clusters where the samples share similar semantics. In the second step, we aggregate the fine-grained clusters based on each cluster’s centroid to establish coarse-grained clustering, facilitating specialization among data experts across different organs.
@@
, where $S=\bigcup S_{i}$ - set of learned cluster centers.
@@
To determine the optimal number of clusters for both steps, we performed a within-cluster sum of squares (WCSS) analysis and used the elbow method to identify the point where adding more clusters provided diminishing returns. This ensured that our clustering approach captured meaningful biological variation while avoiding excessive fragmentation of the data. More details in C.
Figure 2: Two-step clustering framework for data expert construction. Each WSI is divided into patches and encoded into patch-level features. Step 1: Within each tissue type, K-Means clustering is applied to obtain tissue-specific centroids. Step 2: The centroids across tissue types are clustered again to form a final unified set of $K$ clusters (data experts) used for data experts training.
@@ 3.3 Data Experts Training
In this work, we propose to train a set of contrastive models $\{f(\cdot|c)\}$ independently given clusters $C$, such that $c\in C$. The training data for a data expert $\{f(\cdot|c)\}$ is :
@@ 3.4 Slide-level Representation
The proposed expert-based approach provides multiple models to select from during inference rather than relying solely on a single contrastive model. We evaluate three distinct weighting techniques for routing data experts by formulating their outputs as a weighted sum for simplicity:
@@
where $d_{c^{\prime}}=(||h_{v,s}-h_{v,c^{\prime}})||^{2}_{2}$. In the second approach, the outputs from all experts are averaged for each data sample:
@@
In the third approach, the embedding is compared to all centroids to compute distances $\{d_{c}^{\prime}\}_{c^{\prime}\in C}$, and the outputs are combined using a softmax-based weighting:
$p_{3}(c|(h_{v,s},h_{v,c}))=\frac{\exp(-d_{c^{\prime}})}{\sum_{i\in C}\exp(-d_{i})}.$ (9)
After features are obtained for each patch through one of the weighting techniques, a slide-level feature vector can be derived through a learnable aggregation method. Specifically, we employed Attention-Based Multiple Instance Learning (ABMIL) [21], which allows for aggregating patch-level embeddings into a slide-level representation. In this setup, each patch embedding is treated as an instance, and the model learns to assign attention weights to each instance based on its relevance to the slide-level prediction task. ABMIL leverages an attention mechanism to weight each patch’s contribution, capturing spatial heterogeneity across the slide. The resulting slide-level feature vector can then be utilized for various downstream tasks, as described in Section 4.1.
@@ 4.1 Datasets
Each dataset was processed in a similar manner with a fixed stain normalization and $\times224$ patching at magnification 20x. The Visium gene expression data was processed by normalizing each spot to its total count, followed by log normalization. Subsequently, highly variable genes were identified using the Scanpy package [22]. The union of the top 50 most highly variable genes from each sample was used for training and prediction, resulting in 7,986 genes following the method outlined in [17]. Additional details can be found in Appendix.
@@
Subtyping We evaluated the proposed approach on six different subtyping tasks: Non-Small Cell Lung Carcinoma (NSCLC) subtyping on PLCO [23] and CPTAC [24] (two classes), ISUP grading based on Prostate cancer grade assessment (PANDA) challenge [25] (six classes), classification of breast cancer metastases in lymph nodes (Camelyon16) [26, 27] (three classes), ovarian cancer subtypes classification (UBC-OCEAN) challenge [28, 29] (five classes), and PLCO Breast cancer subtyping classification (three classes) [23]. For each dataset where patient identifiers were provided, all patient WSIs were aggregated into one set of vectors for input to ABMIL.
Survival We evaluated the proposed method on five different survival tasks: Breast Invasive Carcinoma (BRCA) from TCGA and PLCO, Non-Small Cell Lung Carcinoma (NSCLC) from PLCO, Uterine Corpus Endometrial Carcinoma (UCEC) from TCGA, Prostate cancer progression-free survival (PRAD) from TCGA.
Prostate Cancer Biochemical Recurrence We evaluated the proposed solution on one prostate cancer biochemical recurrence dataset of whole-mount radical prostatectomies obtained from our instutition.
Gene Biomarker Prediction Following the dataset curation of Kather et al. [30], we evaluated our method on six gene-alteration predictions tasks. For colorectal cancer, we combined TCGA-COAD and TCGA-READ (TCGA-CRC) to predict occurence of driver mutations in KRAS, BRAF, and TP53. For breast cancer, we used TCGA-BRCA to predict receptor status (HER2, PR, ER).